MONT 104N – Modeling the Environment First Writing Assignment, in Two Stages August 21, 2017

The first of the larger paper assignments this semester will deal with one of the ways some of the themes from Frankenstein have been used to attempt to influence public opinion about a contentious topic in the application of science to agriculture—the development of genetically modified organisms (GMOs) such as corn plants that are bred for resistance to pesticides or herbicides, or rice that is genetically modified to include nutrients not found in standard varieties. These have been described by some oponents as “Frankenfoods.” There is certainly an emotional reaction produced by that characterization for lots of people and that, combined with a seeming aversion that many people share to the whole idea of humans intentionally altering genetic material of other organisms, has led to strong resistance to the use of GMOs in human food chains on the part of many people in the U.S. and Europe. On the other hand, there is a pretty strong consensus among scientists that there is no demonstrable difference between the nutrients we would get from food made from these sources and those we get from sources that have not been modified in those ways.

This assignment will have two parts or stages, along the lines of the “They Say, I Say” paradigm:

• Stage 1—The “They Say” part: You will write about two double-spaced pages to paraphrase and summarize the arguments presented in two opinion pieces originally published online and posted on the course homepage:

(a) Frankenfoods: Good for Big Business, Bad for the Rest of Us by Tom Laskawy,

    from grist.org

(b) GMOs Are Not “Frankenfoods,” by Dov Greenbaum and Mark Gerstein, from www.the-scientist.com (this one makes the connection with Mary Shelley’s novel explicitly).

Don’t add your own opinion or take sides between these two points of view yet. This first part of the assignment is geared toward seeing how well you understand and how faithfully you can present the ideas in these two sources. The first stage will be due by email to jlittle@holycross.edu no later than 5:00pm on Friday, September 7.

• I will return the Stage 1 the following week by Wednesday September 12 with comments.

• Stage 2—The “I Say” part: Complete an essay of three to five double-spaced pages in all, using your Stage 1, revised as needed, followed by your own take on these issues. Based on the arguments presented here, what approach seems most reasonable to you? Does the difficulty of feeding the almost 8 billion people alive now justify the use of GMOs? Have biotech/agribusiness companies like Monsanto demonstrated that GMOs are helpful? Do you worry about the potential effects of GMOs in human food sources, or is this just alarmism? Note: If you feel both sides make some convincing points, but you do not have enough information to make a decision or have an opinion either way, it’s OK to say that too. If that is the way you come down, what additional information would you need? Or is this sort of question the kind of thing “ordinary people” should have a voice in deciding at all? Note: There is no one “right answer” here. Reasonable people can come to completely different conclusions, and what this part of the assignment is about is getting you to think through a position of your own and articulate it convincingly. The second stage will be due by email to jlittle@holycross.edu no later than 5:00pm on Monday, September 17.

Guidelines for this assignment

Follow these specifications:

1. Stage 1 is all about conveying what other people have said about why GMOs are so controversial and what we should be doing about them. Don’t prematurely insert your own opinions and try not to let whatever opinions you have about these issues lead you into misrepresentations. Your presentation of the points of view presented in the two pieces should be clear, complete, and it should correctly convey what those authors were trying to say.

2. If you use direct quotations from either opinion piece, identify which one you are drawing from with a parenthetical reference. Here’s an example of what I’m looking for:

Mary Shelley’s Frankenstein has had a major impact by articulating the public’s “visceral fears of scientific and technological innovation” (Greenbaum and Gerstein).

3. If you consult any other sources—this means other books, web sites, etc.—and make use of direct quotations or ideas presented there but not in the assigned pieces, provide enough information for an interested reader to find and check the source you are using. For books, give the author, title, publisher, year of publication, and page where your information is coming from. For web sites, give the full URL, an author if you can identify one, and the date you consulted the site.

4. You can start out with a few sentences about the general background on GMOs, on their labeling as “Frankenfoods,” how that term relates to Mary Shelley’s Frankenstein, and how the general question of whether using GMOs is justified. But don’t spend more than one paragraph doing that. Get to the points made in the two opinion pieces and get specific. When you get to Stage 2, the paper as a whole should be well-organized and “flow” from one point to the next. Do not repeat points unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

5. Your sentences should be clear, readable, and not overly wordy. You should aim for a lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

6. Errors in mechanics (spelling, punctuation, grammar) and usage should not annoy the reader (i.e. Prof. Little). (A word of warning: Your reader tends to be easily annoyed by this sort of thing!) Try as much as possible to avoid those mistakes.